A solenoid is 0.30 m long with 103 turns per meter and carries a current of 5.0 A. what is

the magnitude of the magnetic field at the center of this solenoid?


1
Expert's answer
2021-05-20T20:15:15-0400

we know that magnetic field at center of solenoid is given by - "B= {\mu_0 NI}"


N = "\dfrac {n}{l}" now putting the value of N in above formula , we get


"B=\dfrac {\mu_0 nI}{L}"


"{\mu_0}=4{\pi}\times 10^{-7}" H/m


"n=" 103


"L=" 0.30 m


"I=5.0 A"


now putting all the above equation ,


"B=\dfrac{4{\pi}\times 10^{-7}\times 103 \times 5.0}{0.30}"



"B=2.156\times 10^{-3}" T